Project

General

Profile

NDOS Commissioning Meeting 2011/07/28

Attendance: Jon, Mark Mathis, Minerba, Mat, Martin, Susan, Nick, Dominick, Zeb, Tomas, Jiri, Peter, Athans, Carl B., Art, Andrew

Agenda:

  1. Beam update
  2. Detector update, including:
    1. Beam trigger timing issues
    2. APD recovery status
    3. APD reinstallation plans
  3. DataCheck Update
  4. DAQ Update
  5. DCS Update
  6. AOB

Beam Update

  • NuMI beam came back last night in anti-neutrino mode. We will continue to collect data in this mode until the next (new) NuMI target is available for installation (September). This was a decision made by the MINOS collaboration.
  • We got a couple of hours of beam before a power outage at the lab around midnight.
  • Currently the entire lab is being powered by a single ComEd station (normally there are two). ComEd now understands what repairs are needed for the other station and are working on that. Repairs should finish sometime tonight, and so the lab is planning a site-wide shutdown at 7 am tomorrow so the switch from 1 to 2 stations can be made. Power should be back by 7:30 am.
  • This means that we'll need to stop whatever run we have going around 6:45 am, and power-off as much of the DAQ as possible. Jon will ask Denis or Andrew to shut down the DAQ nicely (remotely).
  • If all goes well, we'll get beam back sometime tomorrow.

Detector Update

  • We have collected data on only 1 dcm the past couple of days due to "NoiseFest 2" (DAQ electronics experts on site are investigating the noisy TECC issues using FEBs on the first di-block).
  • No more golden APDs have been lost in the last week
  • APD recovery:
    • Mat: Leon wants to run some tests in the CalTech enviromental chamber with different amounts of pressure applied to the o-rings. No further APD installation at NDOS will happen until these tests are complete.
    • Leon reported to Jon that Bill Gilbert was working on adjusting the APD teststand Labview program to allow for long (eg, 24 hour) tests.
    • Mat returns on 15th of August, we will wait to install APDs until after he gets back
  • Trigger timing issues (Peter):
    • Andrew N has spent some time analyzing the 1Hz signal, and we have found that the drifts we observed are related to the accelerator clock drifts. However, there is still an issue with the TDU "wigging out" which probably did cause us to lose some beam data.
    • Andrew will implement a scheme to beat two TDUs together such that if they are found to agree, then it is highly likely that they are both ok. The DAQ will keep track of this.
    • in the meantime, hardware will be investigated.

DataCheck

DCS Update

  • Athans:
    • not much new, the developers have been busy dealing with DAQ issues
    • some improvements to GUI, which now has FEB temperatures and TECC drive-currents, graphs, etc. Values displayed are raw numbers, need to be scaled to something that a human can understand.

DAQ Update

  • nothing major to report.

AOB